OpenStudy (anonymous):

HELP (4x^2 - 3x - 2) - (2x^2 - 4x - 4) I got 2x^2 - 7, but thats not one of the choices

OpenStudy (anonymous):

how did you get that answer? Maybe we can show you where you go wrong

OpenStudy (anonymous):

@stabar

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

distribute your - to your equation.

OpenStudy (anonymous):

How did you get that answer

OpenStudy (anonymous):

you got the wrong ans, check your arithmetic again

OpenStudy (anonymous):

I combined the like terms

OpenStudy (anonymous):

i got 2x^2 + x + 2

OpenStudy (anonymous):

that's one of the choices how did you get it?

OpenStudy (anonymous):

distribute your negative to ( 2x^2 - 4x - 4) and you get -2x^2 + 4x + 4 4x^2 - 2x^2 = 2x^2 -3x + 4x = x -2 + 4 = 2 2x^2 + x + 2